OpenStudy (anonymous):

Solve |2y - 2| = y. {2/3, 2} {2, 3} {2, 5} {1/3, 2}

OpenStudy (anonymous):

|2y-2|=y 2y-2=\[+\]y 2y-2 =y 2y-y=2 y=2 2y-2 = -y 2y+y=2 3y=2 y = 2?3